# A note on strong similarity and the Connes embedding problem

### v1

@@ Abstract
Consider $u:C\to N$ a unital cb homomorphism with values in N assumed to be a QWEP von Neumann algebra in $B(H)$. Let $v:N^{\prime}\to B(H)$ be the inclusion map. Then $u$ is strongly similar to a *-homomorphism iff the map $u.v:C\otimes N^{\prime}\to B(H)$ (defined by $u.v(a\otimes b)=u(a)v(b)$ extends to a c.b. map $w$ from $C\otimes_{\max}N^{\prime}$ to $B(H)$. Moreover
@@
Clearly $w$ is still a homomorphism, and hence if it is cb it is similar to a *-homomorphism whence the existence of an invertible $T$ on $H$ (with $\|T\|\|T^{-1}\|=\|w\|_{cb}$) such that $x\mapsto Tw(x)T^{-1}$ is a *-homomorphism. But restricting this to $1\otimes N^{\prime}$ we find that $N^{\prime}\ni y\mapsto TyT^{-1}$ is a *-homomorphism which implies by a simple calculation ($TyT^{-1}=(Ty^{*}T^{-1})^{*}$ and hence $T^{*}Ty=yT^{*}T$) that $|T|\in N^{\prime\prime}=N$. Now $T=U|T|$ with $U$ unitary and the map $C\ni x\mapsto U(|T|u(x)|T|^{-1})U^{-1}$ is a *-homomorphism. Since $U$ is unitary $C\ni x\mapsto(|T|u(x)|T|^{-1})$ is a *-homomorphism, proving the strong similarity of $u$.

### v2

@@ Abstract
Consider $u:C\to N$ a unital cb homomorphism with values in a von Neumann algebra N in $B(H)$. Let $v:N^{\prime}\to B(H)$ be the inclusion map. Then $u$ is strongly similar to a *-homomorphism iff the map $u.v:C\otimes N^{\prime}\to B(H)$ (defined by $u.v(a\otimes b)=u(a)v(b)$ extends to a c.b. map $w$ from $C\otimes_{\max}N^{\prime}$ to $B(H)$. Moreover
@@
Assume $w$ c.b. Clearly $w$ is still a homomorphism, and hence if it is c.b. it must be similar to a *-homomorphism whence the existence of an invertible $T$ on $H$ (with $\|T\|\|T^{-1}\|=\|w\|_{cb}$) such that $x\mapsto Tw(x)T^{-1}$ is a *-homomorphism. But restricting this to $1\otimes N^{\prime}$ we find that $N^{\prime}\ni y\mapsto TyT^{-1}$ is a *-homomorphism which implies by a simple calculation ($TyT^{-1}=(Ty^{*}T^{-1})^{*}$ and hence $T^{*}Ty=yT^{*}T$) that $|T|\in N^{\prime\prime}=N$. Now $T=U|T|$ with $U$ unitary and the map $C\ni x\mapsto U(|T|u(x)|T|^{-1})U^{-1}$ is a *-homomorphism. Since $U$ is unitary $C\ni x\mapsto(|T|u(x)|T|^{-1})$ is a *-homomorphism, proving the strong similarity of $u$.